OpenStudy (themathasker):

WILL FAN AND MEDAL Tell whether the given number is a solution of each equation. 5b + 1 = 16; -3

OpenStudy (tgstudios):

no, because 5(-3) = -15... -15 + 1 = -14... b would need to equal positive 3

OpenStudy (mathmale):

If -3 is a "solution," then "5b + 1 = 16" would be true. If you substitute -3 for b and 5b + 1 = 16 is not true, then b=-3 is NOT a solution.